OpenStudy (anonymous):

Match the number of sides in a polygon on the left with the sum of the angle measures for that polygon on the right. Match Term Definition 9 A) one thousand six hundred twenty degrees 11 B) three thousand four hundred twenty degrees 16 C) two thousand five hundred twenty degrees 21 D) one thousand two hundred sixty degrees

OpenStudy (anonymous):

Yeah lol

OpenStudy (perl):

use the formula sum of angles = 180 * (n-2)

OpenStudy (anonymous):

OK thanks I'll do it now

OpenStudy (perl):

9gon : 180 (9-2) = 1260 11gon: 180 ( 11 - 2) = 1620 16gon: 180( 16-2) = 2520 21gon: 180 (21-2)= 3420

OpenStudy (anonymous):

Thanks I just did It myself using the formula